OpenStudy (anonymous):

Give the required elements of the hyperbola. 2y^2 - 3x^2 = 6 The value of a is: √2 √3 2 3

OpenStudy (anonymous):

Do I have to simplify first, or do I take the bigger number, which is 3?

OpenStudy (anonymous):

to get this into standard form first divide both sides by 6

OpenStudy (anonymous):

(1/3)y^2-(1/2)x^2=1

OpenStudy (anonymous):

agreed. now compare this with the standard for of a hyperbola (centered at the origin):\[\frac{y^2}{b^2}-\frac{x^2}{a^2}=1\]

OpenStudy (anonymous):

OOOOOh , So it's 3?

OpenStudy (anonymous):

a is whichever number is bigger, right?

OpenStudy (anonymous):

a^2 is the number under the x^2

OpenStudy (anonymous):

Hahah thanks. It's so simple, but i'm completely clueless.

OpenStudy (anonymous):

yeah. make sure you realize that a does not equal 2. a^2=2

OpenStudy (anonymous):

@eseidl I'm pretty sure that \(a^2\) is always under the positive one in a hyperbola. \[\frac{x^2}{a^2} - \frac{y^2}{b^{2}} = 1\]\[\frac{y^2}{a^2} - \frac{x^2}{b^2} = 1\]

OpenStudy (anonymous):

GOD flutterING DAMN IT

OpenStudy (anonymous):

@Calcmathlete let me check my text, you may be correct.

OpenStudy (anonymous):

it's still 3 though?

OpenStudy (anonymous):

If I am correct, then \(a^2 = 3\). You ned what \(a\) is.

OpenStudy (anonymous):

square root of 3

OpenStudy (anonymous):

@Calcmathlete I stand corrected. so a^2=3.

OpenStudy (anonymous):

@idontgetconicsections THat works :) \(a = \sqrt{3}\).

OpenStudy (anonymous):

So this conic has foci:\[(0,\pm c)\] where c^2=a^2+b^2=5 c=sqrt 5 It has vertices \[(0,\pm a)\]where a=sqrt 3 and asymptotes \[y=\pm (\frac{a}{b})x\]

OpenStudy (anonymous):

also, It opens up/down